OpenStudy (anonymous):

Solve the equation= 3x[cubed]+18x[squared]-21x=0

OpenStudy (anonymous):

factor out the GCF in this case 3x

OpenStudy (anonymous):

so like 3x[x(squared)+6x-7]=0?

OpenStudy (anonymous):

yes then use product sum or double bubble, whatever you call it on the last part the x^2+6x-7=0

OpenStudy (anonymous):

okay, i understand that but nowi have an x that is squared and one thats not so theyre not like values, meaning i can not combine them, so what do i do?

OpenStudy (anonymous):

you leave the 3x separate here look at this 3x^3+18x^2-21=0 3x(x^2+6x-7)=0 3x(x+7)(x-1)=0 3x=0 x=0 x+7=0 x=-7 x-1=0 x=1 x={-7,0,1}

OpenStudy (anonymous):

ohhh i see! so i can have more than one answer for x, thats allowed?

OpenStudy (anonymous):

heres a trick, whatever the highest exponent in the equation is, equals the number of values you can have example: x^4+x^2=0 has four different values for x

OpenStudy (anonymous):

in this equation the highest is x^3 so you have three values

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem =)